OpenStudy (bbb911):

A pair of crutches needs adjustment. One crutch is 43 inches and the other 41 5/8 inches. Find how much the short crutch should be lengthened to make both crutches the same length.

OpenStudy (ghazi):

simplify 41 5/8 first

OpenStudy (ghazi):

\[41\frac{ 5 }{ 8 }=\frac{ 41*8+5 }{ 8 }=\frac{ 328+5 }{ 8 }=\frac{ 333 }{ 8 }\]

OpenStudy (ghazi):

and then \[43-\frac{ 333 }{ 8 }\]

OpenStudy (bbb911):

would be 333

OpenStudy (ghazi):

it would be i guess 1/8

OpenStudy (bbb911):

then u would have minus them right to get 1/8?

OpenStudy (ghazi):

yea

OpenStudy (bbb911):

okay i did that and got 290/8 now do i simplify that?

OpenStudy (anonymous):

1 3/8

OpenStudy (bbb911):

okay im confused is it 1/8 or 1 3/8

OpenStudy (bbb911):

wouldn't it be 36 1/4?

OpenStudy (anonymous):

It needs 1 3/8 inches (meaning 1.375) to make both crutches equal.

OpenStudy (bbb911):

but how would i get that answer?

OpenStudy (anonymous):

First 5/8 = .625, right? So 41 5/8 = 41.625. Then 43 - 41.625 = 1.375. Now you need to convert it to a fraction -- 1 375/1000 = 1 3/8. Hope it helps? :\

OpenStudy (bbb911):

thank you :)